Suppose that the Sales in December 2019 were 46,595. The sales are expected to grow by 6% in January 2020, by 9% in February 2020, and by 6% in March 2020. After that, the sales are expected to grow by 10%.

The company's cash collection is as follows: 65% is collected in the same month and the remainder in the next month.

What is the cash collection for July, 2020?

Enter your answer rounded off to two decimal points. Do not enter $ or comma in the answer box.
Answer #1

Answer :-

The Sales in December 2019 = $ 46,595

It given ,sales are expected to grow by 6% in January 2020

Sales in January 2020 = $46,595 + ($46,595 × 6%)

=$46595 + $2796 =$ 49391

Sales are grow by 9% in February

Sales in February 2020 = $ 49391 + ($49391 × 9% )

=$ 49391 + $4445 = $53836

sales are grow by 6% in March 2020

Sales in March 2020 = $ 53836 + ($53836 × 6 % )

= $ 53836 + $ 3230 = $57066

After that, the sales are expected to grow by 10%.

Sales in April 2020 = $ 57066 + ($ 57066 × 10 % )

= $ 57066 + $ 5707 = $ 62773

Sales in May 2020 = $ 62773 +($ 62773 × 10% )

= $ 62773 + $ 6277 = $ 69050

Sales in June 2020 = $ 69050 + ( 69050 × 10%)

= $ 69050 +$ 6905 = $ 75955

Sales in July 2020 =$ 75955 + ( $ 75955 + 10% )

= $75955 + $ 7596 = $ 83551

65% of the amount is collected in the May and the remainder in the next month

65% of Amount collected in same month June = 75955 × 65% = $ 49371

Remaining 35% of the amount will collect in the next month ie.July = $ 75955 × 35% = $ 26584

Amount collected in July =( $83551 × 65% ) + 35% of the amount of previous month June

Amount collected in July = $ 54308 +$ 26584

Amount collected in July 2020 = 80892